### v3

@@ 2.2.2 Model Architecture
Similar to textual LLMs, the transformer (Vaswani et al., 2017) is the core model architecture used in recent TSLMs. [^11] Additionally, there are non-tranformer-based architectures such as multi-layer perceptrons (MLP) used in some TSLMs. Of the various models we considered, all of them use a transformer architecture, except for TTM which uses a traditional MLP-based architecture. Transformer-based models leverage self-attention mechanisms to capture long-range dependencies within time series, offering significant advantages in handling sequential data. Self-attention refers to the ability of transformers to focus on relevant parts of the time series data. In a transformer model, each token is projected into a higher dimensional space, allowing the model to capture richer information within that segment. Then, the self-attention layer applies linear transformations and uses the dot product between transformed vectors to identify relevant parts of the time series for each token. This helps in identifying long-range dependencies and relationships, agnostic of the token positions in the input time series. The original transformer architecture utilized two components, namely (i) an encoder, responsible for extracting the relevant information from the input time series and projecting/embedding it into a vector representation and (ii) a decoder, responsible for autoregressive prediction of next token conditioned on observed input sequence of tokens, and encoder output. Recently, several transformer architectures have evolved, notably encoder-decoder, decoder-only, and encoder-only. In the encoder-decoder setup (e.g. Chronos), the encoder maps an input sentence of some language to a continuous representation, and the decoder generates the translation token-by-token using the input representation and previously decoded tokens. In the decoder-only case (e.g. TimesFM, Das et al., 2024), the model only attends to tokens up to the current token, while in the encoder-only case (e.g. Moirai, Woo et al., 2024), the model uses information from the entire sequence to generate tokens. Note that encoder-decoder and decoder-only generate output tokens in an autoregressive manner in contrast to encoder-only architecture.
@@ 2.3 Relationship with Deep Learning
Before moving on to describe the existing TSLMs, it is worth briefly discussing the difference between Deep Learning (DL) models (see for example Torres et al. (2021), Lim and Zohren (2021) and Wen et al. (2022)) and TSLMs. DL models, share many similarities with traditional time-series methods, requiring problem specific training data to learn model parameters and make forecasts. However, they differ the traditional methods in a number of ways – using non-linear architectures, large numbers of learnable parameters, and requiring much larger training data than traditional time series methods. TSLMs, on the other hand, are pretrained on large number of time series datasets. As with FMs, this pretraining allows them to perform “zero-shot” forecasting. This plug-and-play approach makes TSLMs particularly useful and powerful. Furthermore, pretraining allows TSLMs to require much less domain-specific knowledge data compared to deep learning models, and if needed can be fine-tuned on specific datasets.
2.4 State-of-the-art TSFMs
We now briefly describe the TSFM models considered in this paper. [^12] In addition to the models listed in this section, we have also experimented with MOMENT (Goswami et al., 2024), a family of time series transformer models that is pretrained on masked time series prediction task using patching. However this model is still in its very early stages and we have found its performance to be lagging significantly below the other TSFMs we have considered. We therefore did not include it in the comparison below.
@@ 2.4 State-of-the-art TSFMs
(Woo et al., 2024) Salesforce’s Moirai is pretrained on $27$B observations from time series datasets across nine domains. Moirai uses multiple patch (window) size projection layers to capture temporal patterns across various frequencies. It comes in 3 different model sizes – largest of which has about $311$M parameters, which is significant in size compared to earlier models.
@@
(Ekambaram et al., 2024) IBM’s TTM (Tiny Time Mixer), as the name suggests, is a small TSFM with $\leq 1$M parameters that is pretrained on the Monash time series repository Godahewa et al. (2021) and LibCity Jiang et al. (2023) with $\sim 1$B data points. TTM is based on TSMixer architecture (Ekambaram et al., 2023) which uses MLP-based (multi-layer perception) mixing across time steps and features. This enables multivariate forecasting through efficient extraction of temporal and cross-variate patterns for time series forecasting. TTM first normalizes each sequence to have zero mean and unit standard deviation. Then it considers patches (windows) of varying length and varying resolution, since each dataset may have different optimal context to consider. TTM also applies downsampling on high resolution time series (e.g. seconds, minutes) to augment the training dataset.
@@
(Das et al., 2024) Google’s TimesFM is a $200$M parameter TSFM. It is trained on largest corpus of 100 billion time points that includes both real and synthetic time series. TimesFM breaks time series in patches and learns to predict subsequent patches, and uses a standard normalization at patch level.
@@ 3 Econometric Models
Previous work focusing on macroeconomic forecasting has evidenced that models which perform particularly well are Bayesian VARs (BVAR) and Factor Models. See e.g. Stock and Watson (2006), Banbura et al. (2010), Carriero et al. (2016), Carriero et al. (2019), Korobilis and Pettenuzzo (2019), Koop et al. (2019), Chan (2022). We therefore focus on these models, and in this section describe the specific choices we made in terms of their specifications. Starting with the BVARs, we consider both the popular natural conjugate Normal-Inverse Wishart prior, and an alternative specification which preserves conjugacy but allows for cross-shrinkage, a feature that has been shown to be particularly helpful in forecasting. We provide more details below.
3.1 Bayesian VAR with Natural Conjugate Priors
Our starting point is a BVAR with the natural conjugate Normal-inverted Wishart (N-IW) prior. This prior dates back to Zellner (1971) and was later studied by Kadiyala and Karlsson (1997), Kadiyala and Karlsson (1993). In a seminal paper, Banbura et al. (2010) show that it can be successfully applied to a very large cross section of macroeconomic data. Several contributions followed that used this model and prior to handle large macroeconomic datasets.
@@ 3.1 Bayesian VAR with Natural Conjugate Priors
where $t=1,\dots,T$ denotes time. Note that with $p$ lags and $N$ variables, each equation has $K=Np+1$ regressors. By grouping the coefficient matrices in the $N\times$ $K$ matrix $\Phi^{{}^{\prime}}=[\Phi_{c}\ \Phi_{1}\ ...\ \Phi_{p}]$ and defining $x_{t}=(1\ y_{t-1}^{\prime}\ ...\ y_{t-p}^{\prime})^{\prime}$as a vector containing an intercept and $p$ lags of $y_{t}$, the $VAR$ can be written as: $y_{t}=\Phi^{\prime}x_{t}+\varepsilon_{t}.$ Next, after stacking the $t=1,\dots,T$ equations by column and transposing we can rewrite the $VAR(p)$ using more compact matrix notation:
@@
where $Y=$ $[y_{1},..,y_{T}]^{\prime}$, $X=$ $[x_{1},..,x_{T}]^{\prime}$, and $E=[\varepsilon_{1},..,\varepsilon_{T}]^{\prime}$ are $T\times N$, $\ T\times K$and $T\times N$ matrices.
@@
Importantly, the Kronecker product implies that the error variance $\Sigma$ enters the prior variance of the coefficients in a way that is symmetric and proportional to each equation. This in turn implies that the joint prior distribution of $p(\Phi\ ,\Sigma\ )$ is a matricvariate t distribution. The conditional posterior distribution of this model is also N-IW (Zellner (1971)):
@@
where $\bar{\Phi}=\bar{\Omega}(\Omega_{0}^{-1}\Phi_{0}+X^{\prime}Y)$, $\bar{\Omega}=(\Omega_{0}^{-1}+X^{\prime}X)^{-1}$, $\bar{S}=S_{0}+Y^{\prime}Y+\Phi_{0}^{\prime}\Omega_{0}^{-1}\Phi_{0}-\bar{\Phi}^{\prime}\bar{\Omega}^{-1}\bar{\Phi}$, and $\ \bar{v}=v_{0}+T$. We elicit $\Phi_{0},\Omega_{0},S_{0},v_{0}$ as in the Minnesota tradition, and more specific details on this can be found in Appendix A.1. To these priors we add the "sum of coefficients" and "single unit root" priors of Doan et al. (1984) and Sims (1993), which are both discussed also in Sims and Zha (1998). These priors were motivated by the need to avoid having an unreasonably large share of the sample period variation in the data accounted for by deterministic components. Again, more details can be found in Appendix A.1.
Two points worth discussing. First, all of these priors ultimately depend on a small set of hyperparemeters, which we estimate by maximizing the marginal likelihood of the model, and which is available in closed form. For more details see Appendix A.1. Second, note that the posterior variance of the coefficients has a Kronecker structure and therefore the joint prior distribution of $p(\Phi\ ,\Sigma\ |Y)$ is matricvariate-t (and hence the prior is shown to be conjugate). The advantages of this setup are the fact that i) a closed form solution exists for the marginal likelihood, ii) since we know the form of the joint posterior distribution simulations can be performed via simple Monte Carlo sampling (as opposed to e.g. MCMC or SMC), and iii) the Kronecker structure ensures that the operations required to perform such sampling are at most of order $O(N^{3})$.   See Carriero et al. (2015) for a discussion. For these reasons, as shown by Banbura et al. (2010), the BVAR with conjugate prior can handle very large datasets with ease.
3.2 Bayesian VAR with Asymmetric Conjugate Priors
One important limitation of the natural conjugate prior described above is that the Kronecker structure of the prior variance of the coefficients implies that the prior must be symmetric across equations. This in turn rules out cross-variable shrinkage, i.e. the possibility to shrink coefficients on lags of other variables more aggressively than those on own lags. For this reason, Chan (2022) developed an alternative specification that allows for asymmetry in the prior, while maintaining conjugacy. Rewrite the BVAR in (2) in its structural form, i.e.
$Ay_{t}=b+B_{1}y_{t-1}+B_{2}y_{t-2}+\ldots+B_{p}y_{t-p}+u_{t};~{}~{}u_{t}\sim i.i.d.~{}N\left(0,D\right)$ (6)
where $A$ is a lower triangular matrix with ones on the diagonal and $D$ is diagonal. This, in turns, allows for estimation of the model recursively, one equation at a time. It is easy to see that the structural parameters in (6) can be mapped to the reduced-form parameters in (2) by the following transformations: $\Phi_{c}=A^{-1}b$, $\Phi_{j}=A^{-1}B_{j}$ ($j=1,...,p$), and $\Sigma=A^{-1}D\left(A^{-1}\right)^{\prime}$.
Rewrite the $i$-th equation of the BVAR in (6) as
$y_{i,t}=\widetilde{w}_{i,t}\alpha_{i}+\widetilde{x}_{i,t}\beta_{i}+u_{i,t},~{}~{}~{}u_{it}\sim i.i.d.N(0,\sigma_{i}^{2})$ (7)
where $\widetilde{w}_{i,t}=\left(-y_{1,t},\ldots,-y_{i-1,t}\right)$, $\widetilde{x}_{i,t}=\left(1,y_{t-1}^{\prime},\ldots,y_{t-p}^{\prime}\right)$, $\alpha_{i}$ includes the first $i-1$ elements on the $i$-th row of the matrix $A$,  and $\beta_{i}$ includes all the elements from the $i$-th of the matrix obtained after stacking together all the BVAR coefficients as follows: $\left(b,B_{1},\ldots,B_{p}\right)$.
The asymmetric conjugate prior assumes that all the parameters of the BVAR are a priori independent across equations, $p\left(\alpha_{1},\beta_{1},\sigma_{1}^{2},\ldots,\alpha_{N},\beta_{N},\sigma_{N}^{2}\right)=\prod_{i=1}^{N}p\left(\alpha_{i},\beta_{i},\sigma_{i}^{2}\right)$. The priors on the coefficients of each equation $i$ take the form:
$\sigma_{i}^{2}$ $\sim IG\left(\frac{v_{0}+i-N}{2},\frac{s_{i}^{2}}{2}\right)~{}~{}~{}i=1,...,N$ (8)
$\left.\alpha_{i,j}\right|\sigma_{i}^{2}$ $\sim N\left(0,\frac{\sigma_{i}^{2}}{s_{j}^{2}}\right)~{}~{}~{}~{}1\leq j\leq i,i=2,...,N$ (9)
$\left.\beta_{i}\right|\sigma_{i}^{2}$ $\sim N\left(m_{i},\sigma_{i}^{2}V_{i}\right)~{}~{}~{}i=1,...,N.$ (10)
@@ 3.3 Factor Model
Note that this approach (known as the "direct approach") is directly minimizing the $h$-step ahead forecast error, and it requires to perfrom a different regression for each forecast horizon.
@@ 3.4 Bayesian Additive Regression Trees
VARs and factor models assume linearity between endogenous variables and their lags or lags of their linear combinations. Although these models have proven quite successful at forecasting macroeconomic time series, the linearity assumption could be overly restrictive and this could have a negative impact on forecast accuracy. To test this hypothesis, we include Bayesian Additive Regression Trees (BART) in the pool of candidate models that we consider. BART (Chipman et al., 2010) models the conditional mean of the regression model by summing over a large number of trees which are, by themselves, constrained through a regularization prior. The main intuition behind the approach is that while each of these simple trees may only be able to explain a small fraction of the variation in the response variable, summing over a large number of them can describe the data extremely well. Prüser (2019), Huber and Rossini (2022), Huber et al. (2023) and Clark et al. (2023) are recent examples of successful applications of BART with macroeconomic data.
@@
where $X_{-i,t-h}$ denotes all series at time $t-h$ other than $y_{i,t-h}$ and $g\left(X_{-i,t-h}\right)$ is a non-linear function obtained by summing over a larger number of regression trees,
@@
where $g\left(X_{-i,t-h}|\mathcal{T}_{j},m_{j}\right)$ identifies a single tree model with $\mathcal{T}_{j}$ denoting the tree structure and $m_{j}$ is the vector of terminal node parameters associated with $\mathcal{T}_{j}$. Following, Chipman et al. (2010), we set $M=250$ in our application, and specify a prior on the tree structure along the lines they suggested. The $h$-step ahead forecast is then computed as
$\hat{y}_{i,t+h}=\hat{\alpha}_{h}+\hat{\gamma}_{h}y_{i,t}+\widehat{f}\left(X_{-i,t}\right)$ (15)
@@
Abbreviation Description Transformation
PAYEMS All Employees: Total nonfarm 5
INDPRO IP Index 5
FEDFUNDS Effective Federal Funds Rate 1
UNRATE Civilian Unemployment Rate 1
RPI Real personal income 5
DPCERA3M086SBEA Real PCE 5
CMRMTSPLx Real Manu. and TradeIndustries Sales 5
CUMFNS Capacity Utilization: Manufacturing 1
CES0600000007 Avg Weekly Hours: Goods-Producing 1
HOUST Housing Starts, Total 4
S&P 500 S&P’s Common Stock Price Index: Composite 5
T1YFFM 1-Year Treasury C Minus FEDFUNDS 1
T10YFFM 10-Year Treasury C Minus FEDFUNDS 1
BAAFFM Moodys Baa Corporate Bond Minus FEDFUNDS 1
EXUSUKx U.S.-UK Foreign Exchange Rate 5
WPSFD49207 PPI: Final Demand: Finished Goods 5
PPICMM PPI: Metals and metal products 5
PCEPI Personal Consumption Expenditures 5
CES0600000008 Avg Hourly Earnings: Goods-Producing 6
Table 2: Variables in the Medium forecasting model. The column Transformation indicates the transformation of a series $x_{t}$, where: (1) no transformation, (2) $\Delta x_{t}$, (5) $\Delta\log\left(x_{t}\right)$, (6) $\Delta^{2}\log\left(x_{t}\right)$ with $\Delta^{i}$ indicating $i$ th differences.

### v4

@@ 2.2.2 Model Architecture
Similar to textual LLMs, the transformer (Vaswani et al., 2017) is the core model architecture used in recent TSLMs. [^11] Additionally, there are non-tranformer-based architectures such as multi-layer perceptrons (MLP) used in some TSLMs. Of the various models we considered, all of them use a transformer architecture, except for TTM which uses a traditional MLP-based architecture. Transformer-based models leverage self-attention mechanisms to capture long-range dependencies within time series, offering significant advantages in handling sequential data. Self-attention refers to the ability of transformers to focus on relevant parts of the time series data. In a transformer model, each token is projected into a higher dimensional space, allowing the model to capture richer information within that segment. Then, the self-attention layer applies linear transformations and uses the dot product between transformed vectors to identify relevant parts of the time series for each token. This helps in identifying long-range dependencies and relationships, agnostic of the token positions in the input time series. The original transformer architecture utilized two components, namely (i) an encoder, responsible for extracting the relevant information from the input time series and projecting/embedding it into a vector representation and (ii) a decoder, responsible for autoregressive prediction of next token conditioned on observed input sequence of tokens, and encoder output. Recently, several transformer architectures have evolved, notably encoder-decoder, decoder-only, and encoder-only. In the encoder-decoder setup, the encoder maps an input sentence of some language to a continuous representation, and the decoder generates the translation token-by-token using the input representation and previously decoded tokens. In the decoder-only case (e.g. TimesFM, Das et al., 2024), the model only attends to tokens up to the current token, while in the encoder-only case (e.g. Moirai, Woo et al., 2024), the model uses information from the entire sequence to generate tokens. Note that encoder-decoder and decoder-only generate output tokens in an autoregressive manner in contrast to encoder-only architecture.
@@ 2.3 Relationship with Deep Learning
Before moving on to describe the existing TSLMs, it is worth briefly discussing the difference between Deep Learning (DL) models (see for example Torres et al. (2021), Lim and Zohren (2021) and Wen et al. (2022)) and TSLMs. DL models share many similarities with traditional time-series methods, requiring problem specific training data to learn model parameters and make forecasts. However, they differ the traditional methods in a number of ways – using non-linear architectures, large numbers of learnable parameters, and requiring much larger training data than traditional time series methods. TSLMs, on the other hand, are pretrained on large number of time series datasets. As with FMs, this pretraining allows them to perform “zero-shot” forecasting. This plug-and-play approach makes TSLMs particularly useful and powerful. Furthermore, pretraining allows TSLMs to require much less domain-specific knowledge data compared to deep learning models, and if needed can be fine-tuned on specific datasets.
2.4 State-of-the-art TSLMs
We now briefly describe the TSLM models considered in this paper. [^12] In addition to the models listed in this section, we have also experimented with MOMENT (Goswami et al., 2024), a family of time series transformer models that is pretrained on masked time series prediction task using patching. However this model is still in its very early stages and we have found its performance to be lagging significantly below the other TSLMs we have considered. We therefore did not include it in the comparison below.
@@ 2.4 State-of-the-art TSLMs
(Woo et al., 2024) Salesforce’s Moirai is pretrained on 2727B observations from time series datasets across nine domains. Moirai uses multiple patch (window) size projection layers to capture temporal patterns across various frequencies. It comes in 3 different model sizes – largest of which has about 311311M parameters, which is significant in size compared to earlier models.
@@
(Ekambaram et al., 2024) IBM’s TTM (Tiny Time Mixer), as the name suggests, is a small TSLM with $\leq 1$M parameters that is pretrained on the Monash time series repository Godahewa et al. (2021) and LibCity Jiang et al. (2023) with $\sim 1$B data points. TTM is based on TSMixer architecture (Ekambaram et al., 2023) which uses MLP-based (multi-layer perception) mixing across time steps and features. This enables multivariate forecasting through efficient extraction of temporal and cross-variate patterns for time series forecasting. TTM first normalizes each sequence to have zero mean and unit standard deviation. Then it considers patches (windows) of varying length and varying resolution, since each dataset may have different optimal context to consider. TTM also applies downsampling on high resolution time series (e.g. seconds, minutes) to augment the training dataset.
@@
(Das et al., 2024) Google’s TimesFM is a 200200M parameter TSLM. It is trained on largest corpus of 100 billion time points that includes both real and synthetic time series. TimesFM breaks time series in patches and learns to predict subsequent patches, and uses a standard normalization at patch level.
@@ 3 Econometric Models
Previous work focusing on macroeconomic forecasting has evidenced that models which perform particularly well are Bayesian VARs (BVARs) and Factor Models. See e.g. Stock and Watson (2006), Banbura et al. (2010), Carriero et al. (2016), Carriero et al. (2019), Korobilis and Pettenuzzo (2019), Koop et al. (2019), Chan (2022). We therefore start from these models. We also include neural networks and regression trees in the comparison, as recent work by Hyndman and Athanasopoulos (2018), Huber et al. (2023), and Clark et al. (2023) show that in some instances allowing for non-linearities can help improve the accuracy of macroeconomic forecasts. Below we describe the specific choices we made in terms of the model specifications.
3.1 Bayesian VAR
We consider the popular BVARs with natural conjugate Normal-inverted Wishart (N-IW) priors. This prior dates back to Zellner (1971) and was later studied by Kadiyala and Karlsson (1993, 1997). In a seminal paper, Banbura et al. (2010) show that it can be successfully applied to a very large cross section of macroeconomic data. Several contributions followed using this model and prior to handle large macroeconomic datasets.
@@ 3.1 Bayesian VAR
where $t=1,\dots,T$ denotes time. Note that with $p$ lags and $N$ variables, each equation has $K=Np+1$ regressors. By grouping the coefficient matrices in the $N\times$ $K$ matrix $\Phi^{{}^{\prime}}=[\Phi_{c}\ \Phi_{1}\ ...\ \Phi_{p}]$ and defining $x_{t}=(1\ y_{t-1}^{\prime}\ ...\ y_{t-p}^{\prime})^{\prime}\$as a vector containing an intercept and $p$ lags of $y_{t}$, the $VAR$ can be written as: $y_{t}=\Phi^{\prime}x_{t}+\varepsilon_{t}.$ Next, after stacking the $t=1,\dots,T$ equations by column and transposing we can rewrite the $VAR(p)$ using more compact matrix notation:
@@
where $Y=$ $y_{1},..,y_{T}]^{\prime}$, $X=$ $x_{1},..,x_{T}]^{\prime}$, and $E=[\varepsilon_{1},..,\varepsilon_{T}]^{\prime}$ are $T\times N$, $T\times K\$and $T\times N$ matrices.
@@
The conditional posterior distribution of this model is also N-IW:
@@
This model can be used for very large cross sections because the Kronecker product in (4) implies that the error variance $\Sigma$ enters the prior variance of the coefficients in a way that is symmetric and proportional to each equation. This fact, together with the fact that in a VAR every equation has the same regressors, implies that the posterior distribution of $p(\Phi,\Sigma)$ is a matricvariate-t. This presents several advantages as in this case i) a closed form solution exists for the marginal likelihood, ii) simulations can be performed via direct Monte Carlo (MC) sampling (as opposed to e.g. Markov Chain MC or Sequential MC), and iii) the operations required to perform such MC sampling are at most of order $O(N^{3})$.  See Carriero et al. (2015) for a discussion. For these reasons, as shown by Banbura et al. (2010), the BVAR with conjugate prior can handle very large data sets with ease. [^13] It is worth pointing out that one important limitation of the natural conjugate prior described in this section is that the Kronecker structure of the prior variance of the coefficients implies that the prior must be symmetric across equations. This in turn rules out cross-variable shrinkage, i.e. the possibility to shrink coefficients on lags of other variables more aggressively than those on own lags. Chan (2022) developed an alternative specification that allows for asymmetry in the prior, while maintaining conjugacy. We experimented with this alternative and found the results to be broadly similar; therefore, we did not include them in the paper. These results are available from the authors on request.
@@ 3.2 Factor Model
Note that this approach (known as the "direct approach") is directly minimizing the $h$-step ahead forecast error, and it requires to perform a different regression for each forecast horizon.
3.3 Neural Network Autoregression (NNAR)
VARs and factor models assume linearity between endogenous variables and their lags or lags of their linear combinations. Although these models have proven quite successful at forecasting macroeconomic time series, the linearity assumption could be overly restrictive and this could have a negative impact on forecast accuracy. To test this hypothesis, we also include in the comparison the Neural Network Autoregression (NNAR) model of Hyndman and Athanasopoulos (2018) and popularized through the forecast package published by the same authors. The NNAR model combines the structure of classical autoregressive models with the flexibility of feed-forward neural networks. In particular, the model is specified as:
$y_{i,t}=f(y_{i,t-1},y_{i,t-2},\dots,y_{i,t-p})+\varepsilon_{i,t},$
where $f(\cdot)$ is a nonlinear function represented by a single-hidden-layer feed-forward neural network with $k$ hidden nodes, and $\varepsilon_{i,t}$ is a white noise error term. The model is denoted as NNAR(p,k), analogous to an AR(p) model in traditional time series analysis. The NNAR model is based on a single hidden layer neural network trained using backpropagation, where we set the number of lagged inputs to $p=12$. Furthermore, to avoid overfitting, the model uses weight decay (L2 penalty) during training. Forecasts from the NNAR model are generated iteratively, and averaging over multiple networks trained with different initial weights, thereby capturing model uncertainty. Despite its simplicity compared to modern deep learning models, the NNAR model often performs competitively on short-term forecasting tasks, especially for time series with nonlinear dynamics and limited data. [^14] We have also experimented with DeepAR (Salinas et al., 2020), a sigificantly more sophisticated approach leveraging recurrent neural networks that trains the network using all available time series at once, but found it overall to not be more accurate than this simpler alternative.
@@ 3.4 Bayesian Additive Regression Trees
Lastly, we include in the pool of candidate models Bayesian Additive Regression Trees (BART). BART (Chipman et al., 2010) models the conditional mean of the regression model by summing over a large number of trees which are, by themselves, constrained through a regularization prior. The main intuition behind the approach is that while each of these simple trees may only be able to explain a small fraction of the variation in the response variable, summing over a large number of them can describe the data extremely well. Prüser (2019), Huber and Rossini (2022), Huber et al. (2023) and Clark et al. (2023) are recent examples of successful applications of BART with macroeconomic data.
@@
where $X_{-i,t-h}$ denotes all series at time tht-h other than $y_{i,t-h}$ and $g\left(X_{-i,t-h}\right)$ is a non-linear function obtained by summing over a larger number of regression trees,
@@
where $g\left(X_{-i,t-h}|\mathcal{T}_{j},m_{j}\right)$ identifies a single tree model with $\mathcal{T}_{j}$ denoting the tree structure and $m_{j}$ is the vector of terminal node parameters associated with $\mathcal{T}_{j}$. Following Chipman et al. (2010), we set $M=250$ in our application and specify a prior on the tree structure along the lines they suggested. The $h$-step ahead forecast is then computed as
$\widehat{y}_{i,t+h}=\widehat{\alpha}_{h}+\widehat{\gamma}_{h}(L)y_{i,t}+\widehat{f}\left(X_{-i,t}\right)$ (10)